OpenStudy (anonymous):

Write a polynimial function of least degree with integral coefficeints havung zeros that include -1 and 1+2i

OpenStudy (anonymous):

complex roots always occur in pairs ,so zeros are \[-1,1+2 \iota ,1-2 \iota\] eq.of polynomial is \[\left( x+1 \right)\left( x-1+2 \iota \right)\left( x-1- 2 \iota \right)=0\] \[\left( x+1 \right)\left\{ \left( x-1 \right)+2 \iota \right\}\left\{ \left( x-1 \right) -2 \iota\right\}=0\] \[\left( x+1 \right)\left\{ \left( x-1 \right)^2-4 \iota^2 \right\}=0,\iota^2=-1\] simplify and form the equation.